OpenStudy (anonymous):

log(x+5)=log(x)+log(5) solve this equation

OpenStudy (anonymous):

x = 5

OpenStudy (anonymous):

why though?

OpenStudy (anonymous):

log(x-1) + log(5x) = 2 becomes: log((x-1)*(5x)) = 2 this becomes: log(5x^2 - 5x) = 2 the basic rule of logarithms states: y = log(a,x) if and only if a^y = x in your equation, a base of 10 is implied. your equation of log(5x^2 - 5x) = 2 is really: log(10,(5x^2-5x)) = 2 which means: log of (5x^2-5x) to the base of 10 = 2 using the basic rule of logarithms, this equation becomes: log(10,(5x^2-5x)) = 2 if and only if: 10^2 = 5x^2-5x this becomes: 100 = 5x^2 - 5x subtract 100 from both sides of this equation to get: 5x^2 - 5x - 100 = 0 which is a quadratic equation. divide both sides of this equation by 5 to get: x^2 - x - 20 = 0 this factors out to be: (x-5)*(x+4) = 0 solve for x to get: x = 5 or x = -4 substitute in original quadratic equation to confirm these answers are good. 5x^2 - 5x = 100 becomes: 80 + 20 = 100 when x = -4 and becomes: 125 - 25 = 100 when x = 5. both solutions are good. plug these solutions into your original equation that you started with to get: log(x-1) + log(5x) = 2 becomes: log(4) + log(25) = 2 when x = 5. solving this equations gets 2 = 2 confirming x = 5 is a good solution. log(x-1) + log(5x) = 2 becomes: log(-5) + log(-20) = 2 when x = -4 this solutions is not valid because you can't take the log of a negative number. your only valid solution is: x = 5

OpenStudy (anonymous):

ok I could have googled that too...lol